OpenStudy (anonymous):

3x^2+2=-8x <--- solve by the quadratic formula , set the equation to "0" first. must be in simplest radical form, show work please

OpenStudy (dumbcow):

coefficients are a = 3 b = 8 c = 2 substitute these into quadratic formula

OpenStudy (anonymous):

how i get the equation to equal to zero first?

OpenStudy (anonymous):

just add 8x to both sides.

OpenStudy (anonymous):

i got x=\[\pm8\]\[\sqrt{8^2-4(3)(2)}\] /2(3) is that right?

OpenStudy (anonymous):

Should be \[x = \frac{-8 \pm \sqrt{8^2-4(3)(2)}}{2(3)}\] So yeah, the 8 out front is negative, and the plus or minus goes between the -b and the square root. Otherwise that's right.

OpenStudy (anonymous):

so what would i do next?

OpenStudy (anonymous):

Do the multiplication in the square root part and the denominator, then simplify it.

OpenStudy (anonymous):

x=-8 \[\pm\] \[\sqrt{8^2-36}\]/6

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

4(3)(2) = 24\[\implies \sqrt{8^2 - 4(3)(2)} \]\[= \sqrt{8^2 - 24} \]\[= \sqrt{64 - 24}\]\[=\sqrt{40}\]\[=\sqrt{4\cdot 10} \]\[=2\sqrt{10}\]

OpenStudy (anonymous):

so you can replace the square root you have with 2 root(10)

OpenStudy (anonymous):

how you get 24?

OpenStudy (anonymous):

would we put the -8 and that upside down tee looking like before the square root problems you did?

OpenStudy (anonymous):

that upside-down t is called the plus or minus sign. and yes.